# Dissipative spatiotemporal soliton in a driven waveguide laser

### v1

@@ Abstract
A distributed Kerr-lens mode locking regime can be realized in a waveguide laser by spatial profiling of the pump beam, thus creating a spatio-temporal soliton. Additional slow temporal modulation of the pump source stabilizes the spatio-temporal solution in a broad range of parameters, which are defined by the dynamic gain saturation. We choose a Cr:ZnS waveguide laser as a practical example, but such a regime is feasible in various waveguide and fiber oscillators. A far-reaching analogy with Bose-Einstein condensates allows using this approach to stabilization of the weakly dissipative BECs.
journal: jocn
@@ 1 Introduction
The unprecedented progress in the field of mesoscopic physics in its general context, which comprises control of macro-quantum states like Bose-Einstein condensate (BEC) [1, 2, 3, 4] and quantum supersolids [5], plasma and turbulence [6, 7, 8], attosecond [9], relativistic and high-energy physics [10, 11, 12, 13], metaphorical modeling and quantum computations [14, 15, 16, 17], physics of ultra-strongly coupled light-matter systems [18], nano-processing [19, 20, 21], and other ranges of science, medicine, and technology, became feasible due to the growth understanding of strongly nonlinear phenomena in far-from-equilibrium systems [22, 23]. The coherent, self-organized, and well-localized patterns, namely solitons, play a crucial role in most of these phenomena. The class of such structures, dissipative solitons (DSs), is particularly important because DS develops in open systems, where energy exchange with an environment defines its integrity, stability, and coherence [24, 25]. In photonics, DSs are a road to generating ultrafast, robust, and energy-scalable pulses in mode-locked lasers [26, 27] bringing a high-field physics on tabletops of a mid-level university lab [12, 28].
@@
The modern tendencies in generating and exploring laser DSs are based on a controllable enhancement of self-organizing effects induced by nonlinearities and affected both temporal and spatial degrees of freedom [23]. A partial realization of this concept is a distributed Kerr-lens mode-locking (DKLM) [29, 30, 31]. In a solid-state laser, this technique uses a nonlinear medium that affects a laser beam via self-focusing, thereby changing an effective gain in a laser resonator [32]. The transversal spatial structure of a field in this type of device is close to a fundamental mode of laser oscillator and, thus, trivial, so only longitudinal modes need synchronization. A remarkable breakthrough in energy scalability of such oscillators was demonstrated in both anomalous and normal group-delay dispersion regimes (ADR and NDR, respectively) [33, 34, 30]. It was found that such broad-range scalability could be explained by energy out/in-flows induced by the nontrivial phase structure of DS [27, 35], and the underlying theory of the DS resonance (DSR) [36] were developed.
@@
Using DS in the NDR fiber lasers [37] exploits the enhanced and well-controllable nonlinearities. That bridges fiber and solid-state ultrafast laser photonics [23]. The keystone is utilizing a nonlinear propagation of many interacting spatial modes (e.g., in the so-called multimode fibers, MMF) [38, 39]. Such a nonlinear spatial mode coupling could be caused by nonlinear refraction (or attractive boson interaction in BEC [4]) characterized by the coefficient $n_{2}$. A “confining potential” defined by spatial-dependent refractive index $n(x,y)$ (“graded refractive index fibers,” or GRIN, nonlinear lattices [40, 41], or laser-induced confinement in BEC) enhances the relaxation of the higher-order modes to a ground-state, i.e., lowest-order spatial mode.
@@
In this work, we propose utilizing a nonlinear mode condensation for DKML aimed at a spatio-temporal mode-locking (STML) [42, 43].
@@
Such an approach utilizes a dissipative (i.e., $\Im[n(x,y)]\neq 0$, where $n$ is, e.g., a complex index of refraction in optics) trapping potential [44] for a spatiotemporal DS (STDS) without involving other nonlinear dissipative processes. The 2D dissipative potential, unlike that in Refs. [45, 46] transversal $(x,y)$-confinement induced by a complex $n(x,y)$ can be associated with a so-called “cigar-type” trapping potential in a weakly dissipative BEC or GRIN (see Table 1) [15]. Then, an evolutionary (“slow”) coordinate $T$ is associated with time for BEC or propagation distance $Z$ in a fiber or cavity round-trip in a laser.
@@
A further step is to introduce “longitudinal” (“slow” time $T$-dependent, or propagation length $Z$-dependent) modulation of $\Re[n(x,y,T)]$ (or $\Re[n(x,y,Z])$, which is one of the management techniques stabilizing an ST soliton in BEC [47]. In a mode-locked laser, it is also possible to identify a “fast” coordinate (local time $t$) associated with the frame co-moving with a soliton. The dependence of trapping potential on that coordinate allows a 3D confinement corresponding to a “pancake-like” potential in BEC. In a laser, this corresponds to an active mode-locking driven by an external synchronous phase modulation[48, 49, 50], which varies as $\Re[\Delta n(t)]\propto\cos(2\pi t/T_{R})$, where $T_{R}$ is a cavity round-trip period. This modulation is slow (pulse duration $\tau_{p}$ is much shorter than $T_{R}$) and can be described by $\Re[\Delta n(t)]\propto t^{2}+\mathrm{H.O.T.}$. It was found that such a modulation stabilizes a driven cavity DS [51, 52] and STDS [53]. Here, we intend to exploit a concept of 3D dissipative confinement to develop a spatiotemporal mode-locking mechanism to generate stable and energy-scalable STDS (“light bullet”). This technique can be considered as a road to forming a stable mass-scalable STDS in BEC and pattern formation in dissipative turbulent systems, which could be modeled by the dissipative version of the Gross-Pitaevskii equation (GPE) [4, 54]. In contrast to [55, 53], a transversely graded dissipation (i.e., $\Im[n(x,y)]\neq 0$) will be supplemented by dissipative trapping along $t$-coordinate ($\Im[n(t)\neq 0$), which is a result of periodical loss modulation (or synchronous pumping) synchronized with the laser cavity round-trip. For BEC, this corresponds to a “pancake-like” dissipative confining potential.
Table 1: Correspondences between photonics and BEC [15]
Laser
BEC
Propagation coordinate $Z$
$\leftrightarrow$
time $T$
Pulse local time $t$
$\leftrightarrow$
third spatial coordinate $z$
Diffraction + anomalous GDD
$\leftrightarrow$
boson kinetic energy
Kerr-nonlinearity
$\leftrightarrow$
boson attractive colliding potential
GRIN
$\leftrightarrow$
cigar-like ($x$, $y$) trapping potential
Non-selective loss
$\leftrightarrow$
homogeneous BEC dissipation
@@
Graded gain + local time modulation
$\leftrightarrow$
pancake ($x,y,z$)-shaped inflow of bosons
Spectral dissipation
$\Leftrightarrow$
“kinetic cooling”
@@ 2.1 Model
Let us consider a waveguide laser composed of a fiber or crystalline waveguide active medium enclosed in a Fabri-Perót or ring resonator and driven by a pump beam with a parabolic-like transverse $G(x,y)-$ profile and a pump intensity varying with time $t$ at the frequency synchronized with a laser repetition rate. The $t-$coordinate corresponds to a “local time” (in a coordinate frame co-moving with an STDS) in photonics or to a longitudinal space coordinate for BEC (see Fig. 1 and Table 1) [15].
Figure 1: Multi-scale scheme with a local-time coordinate $t$ synchronized with an STDS group velocity and a global evaluation coordinate $Z$ (see Table 1). The gain (filled by a red) is profiled along the radial coordinate $r$ and the local time $t$. The last is synchronized with a laser round-trip period.
Such pump beam acts as a “soft aperture” guiding the laser beam and controlling (“cleaning”) its mode structure so that a laser field $a(z,x,y,t)$ tends to condensate into a lowest-order mode. Simultaneously, $t-$modulation provides confinement in the time domain. Thus, this scheme corresponds to a three-dimensional (“pancake-like”) dissipative confinement for a BEC or a hybrid mode-locking under synchronous pumping or an active assisting mode locking by synchronous loss modulation [56, 57, 50, 52].
For modeling a (3D+1)-dimensional trapped STDS, we use a distributed model describing the evolution of the $(z,x,y,t)$-dependent field $a$ propagating along the $z$-axis (see Table 1 for comparison with BEC) of an active waveguide with the linear and nonlinear refractive indexes $n(x,y)$ and $n_{2}$, respectively. Under the action of $(x,y,t)$-graded pump (Fig. 1), the evolution can be described by a sort of the driven nonlinear Helmholtz equation:
@@
where $\Delta_{x,y}$ is a Laplace operator in the Cartesian coordinates, $n_{0}$ is a waveguide cladding refractive index, $k_{0}$ is a free-space wave number, and the second-order derivative over $z$ contributes only for large beam numerical apertures NA [58, 59]. The terms in Eq. 1 define: (I) – diffraction, (II) – group-delay dispersion with the coefficient $\beta_{2}$, (III) – spectral dissipation with the squared inverse bandwidth $\tau$, (IV) – trapping potential induced by GRIN, (V) – $(x,y,t)$- graded saturated gain with the coefficient $G$ and linear net-loss $L$, and (VI) – Kerr-nonlinearity. One has to emphasize the decisive contribution of a saturable and transversely profiled gain $G(x,y,t)$ in Eq. (1).
For a low NA, one may omit the $\frac{\partial^{2}}{\partial z^{2}}$-term in Eq. (1). That leads to the nonlinear driven Schrödinger, or the Lugiato-Lefever equation (LLE), which is a test bed model for the study of solitons and non-equilibrium patterns in driven Kerr resonators, micro-cavity and fiber lasers, VCSEL, etc. [60, 61, 62, 63].
The rescaling $\psi=a\exp(-iV_{0}z)$ ($V_{0}=(k_{0}/2n_{0})(n_{1}^{2}-n_{0}^{2})$, and $n_{1}$ is a waveguide core refractive index) leads to the dimensionless mean-field LLE [64, 65, 66]:
@@
$\left\{(X^{2}+Y^{2})-\left|\psi\right|^{2}-i\left[\Lambda+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}\right]\right\}\psi,$ (2)
for the normalizations shown in Table 2 [64, 67].
@@
In this work, we assume the parabolic-like transverse profile of the gain beam with the size. Such transversely confined nonlinear gain acts as a “soft aperture” with an effective size $\approx w_{p}\sqrt{G_{0}-\Lambda}$ (Table 2). Our model introduces a gain modulation along the $t-$axis described in a parabolic approximation by the modulation coefficient $\omega$ [50, 57]. The last resembles the mechanism of the so-called active mode-locking (AM) [68]. In contrast to [53], the proposed method does not use additional phase modulation but only a pumping pulse train synchronized with a resonator period. For simplicity, we manipulate with a graded saturated gain $G(X,Y,t)=G(0,0,0)\times\left[1-\kappa(X^{2}+Y^{2})-\omega t^{\prime 2}\right]$, where $G(0,0,0)$ or $G_{0}$ (see Table 2) corresponds to a saturated gain along a waveguide axis at $t^{\prime}=0$, and $\kappa=G_{0}w_{p}^{-2}$, $\omega=G_{0}\mathscr{T}^{-2}$ ($\mathscr{T}$ is a “pump pulse duration”). Such an approximation is valid under the condition of quasi-steady state operation of a laser in the vicinity of the lasing threshold, where the saturated net-loss coefficient $\Lambda^{\prime}=L-G_{0}$ is close to zero [55].
@@
Thus, one may treat the resulting equation (2) as a dissipative extension of the driven Gross-Pitaevskii equation, which is a well-known tool for analyzing trapped BEC (Table 1) [4].
Table 2: Normalizations for the LLE (2)
Value
Normalizations and definitions
$(X,Y)$
transverse coordinates, $(x,y)/w_{T}$
$Z$
propagation distance, $z/\zeta$
$t^{\prime}$
local time, $t/\sqrt{|\beta_{2}|\zeta}$
$\zeta$
propagation scale, $k_{0}n_{o}w_{T}^{2}$
$w_{T}$
transverse scale, $4]{w_{0}^{2}/k_{0}^{2}n_{0}\delta}$
$\delta$
GRIN contrast, $n_{1}-n_{0}$
$\left(k_{0}/2n_{0}\right)\left(n_{1}^{2}-n(x,y)^{2})\right)$
GRIN confining potential,  $\simeq k_{0}\left(n_{1}-n(r)\right)$
$\left|\psi\right|^{2}$
intensity, $k_{0}n_{2}\zeta\left|a\right|^{2}$
$\Lambda^{\prime}$
net-loss, $\Lambda-G_{0}$
$\Lambda$
averaged loss, $LL_{w}/\zeta$
$\kappa$
spatial confinement parameter, $G_{0}/w_{p}^{2}$
$\omega$
temporal confinement parameter, $G_{0}/\mathscr{T}^{2}$
$w_{p}\sqrt{G_{0}-\Lambda}$
soft aperture size
$\tau$
squared inverse spectral filter, bandwidth normalized to $|\beta_{2}|\zeta$
$k_{0}$
wave-number, $2\pi/\lambda$
$\lambda$
central wavelength
$n_{1}$
GRIN core refractive index
$n_{0}$
GRIN cladding refractive index
$n_{2}$
nonlinear refractive index
$L$
loss coefficient
$L_{w}$
waveguide (“lattice”) length
$G_{0}$ or $G(t=0,x=0,y=0)$
maximum saturated gain coefficient
$w_{p}$
pump beam size
$\mathscr{T}$
pump pulse width

### v2

@@ Abstract
A distributed Kerr-lens mode locking regime can be realized in a waveguide laser by spatial profiling of the pump beam, thus creating a spatio-temporal soliton. Additional slow temporal modulation of the pump source stabilizes the spatio-temporal solution in a broad range of parameters defined by the dynamic gain saturation. We choose a Cr:ZnS waveguide laser as a practical example, but such a regime is feasible in various waveguide and fiber oscillators. A far-reaching analogy with Bose-Einstein condensates allows this approach to stabilize the weakly dissipative BECs.
keywords:
waveguide laser , spatiotemporal soliton , mode-locking , mid-infrared lasers
journal: Optics Communications
[…]
[first]organization=Department of Physics, Norwegian University of Science and Technology, addressline=, city=Trondheim, postcode=7491, country=Norway
[…]
[second]organization=Institut für Photonik, TU Wien, addressline=Gusshausstrasse 27/387, city=Vienna, postcode=1040, country=Austria
[…]
[third]organization=Department of Physics, Norwegian University of Science and Technology, addressline=, city=Trondheim, postcode=7491, country=Norway
[…]
[fifth]organization=ATLA lasers AS, addressline=Richard Birkelands vei 2B, city=Trondheim, postcode=7034, country=Norway
@@ 1 Introduction
The unprecedented progress in the field of mesoscopic physics in its general context, which comprises control of macro-quantum states like Bose-Einstein condensate (BEC) Kevrekidis et al. (2008); Luo et al. (2021); Cazalilla et al. (2011); Carretero-González et al. (2008) and quantum supersolids Boninsegni and Prokof’ev (2012), plasma and turbulence Robinson (1997); Dyachenko et al. (1989); Kuznetsov et al. (1986), attosecond Krausz and Ivanov (2009), relativistic and high-energy physics Mourou et al. (2006); England et al. (2014); Südmeyer et al. (2008); Gallerati et al. (2022), metaphorical modeling and quantum computations Faccio (2012); Kalashnikov and Wabnitz (2021); Faccio et al. (2010); Averin et al. (2012), physics of ultra-strongly coupled light-matter systems Forn-Díaz et al. (2019), nano-processing Gattass and Mazur (2008); Andrews et al. (2010); Vorobyev and Guo (2013), and other ranges of science, medicine, and technology, became feasible due to the growth understanding of strongly nonlinear phenomena in far-from-equilibrium systems Cross and Hohenberg (1993); Fu et al. (2018). The coherent, self-organized, and well-localized patterns, namely solitons, play a crucial role in most of these phenomena. The class of such structures, dissipative solitons (DSs), is particularly important because DS develops in open systems, where energy exchange with an environment defines its integrity, stability, and coherence Purwins et al. (2010); Akhmediev and Ankiewicz (2008). In photonics, DSs are a road to generating ultrafast, robust, and energy-scalable pulses in mode-locked lasers Brabec and Krausz (2000); Grelu and Akhmediev (2012) bringing a high-field physics on tabletops of a typical university lab Südmeyer et al. (2008); Brons (2017).
@@
The modern tendencies in generating and exploring laser DSs are based on a controllable enhancement of self-organizing effects induced by nonlinearities and affected by both temporal and spatial degrees of freedom Fu et al. (2018). A partial realization of this concept is a distributed Kerr-lens mode-locking (DKLM) Sorokina et al. (1996); Zhang et al. (2022); Demesh et al. (2023). In a solid-state laser, this technique uses a nonlinear medium that affects a laser beam via self-focusing, thereby changing an effective gain in a laser resonator Brabec et al. (1992). It was demonstrated that one can even completely decouple the role of the nonlinear medium and the gain medium Sorokina et al. (1996). The latter opens up the way to power scaling of ultra-short pulsed lasers, especially in waveguide configuration. The transversal spatial structure of a field in this type of device is close to a fundamental mode of laser oscillator and, thus, trivial, so only longitudinal modes need synchronization. A remarkable breakthrough in energy scalability of such oscillators was demonstrated in both anomalous and normal group-delay dispersion regimes (ADR and NDR, respectively) Baer et al. (2012); Pronin et al. (2012); Zhang et al. (2022). It was found that such broad-range scalability could be explained by energy out/in-flows induced by the nontrivial phase structure of DS Grelu and Akhmediev (2012); Kalashnikov (2012), and the underlying theory of the DS resonance (DSR) Chang et al. (2008) was developed.
@@
Using DS in the NDR fiber lasers Wise et al. (2008) exploits the enhanced and well-controllable nonlinearities. That bridges fiber and solid-state ultrafast laser photonics Fu et al. (2018). The keystone is utilizing a nonlinear propagation of many interacting spatial modes (e.g., in the so-called multimode fibers, MMF) Mondal et al. (2020); Piccardo et al. (2021). Such a nonlinear spatial mode coupling could be caused by nonlinear refraction (or attractive boson interaction in BEC Carretero-González et al. (2008)) characterized by the coefficient $n_{2}$. A “confining potential” defined by spatial-dependent refractive index $n(x,y)$ (“graded refractive index fibers,” or GRIN, nonlinear lattices Matuszewski et al. (2006); Kartashov et al. (2011), or laser-induced confinement in BEC) enhances the relaxation of the higher-order modes to a ground-state, i.e., lowest-order spatial mode.
@@
In this work, we propose utilizing a nonlinear mode condensation for DKML aimed at a spatio-temporal mode-locking (STML) Wright et al. (2017, 2020).
@@
Such an approach utilizes a dissipative (i.e., $\Im[n(x,y)]\neq 0$, where $n$ is, e.g., a complex index of refraction in optics) trapping potential Siegman (2003) for a spatiotemporal DS (STDS) without involving other nonlinear dissipative processes. The 2D dissipative potential, unlike that in Refs. Sun et al. (2023); Li et al. (2022) transversal $(x,y)$-confinement induced by a complex $n(x,y)$ can be associated with a so-called “cigar-type” trapping potential in a weakly dissipative BEC or GRIN (see Table 1) Kalashnikov and Wabnitz (2021). Then, an evolutionary (“slow”) coordinate $T$ is associated with time for BEC or propagation distance $Z$ in a fiber or cavity round-trip in a laser.
@@
A further step is to introduce “longitudinal” (“slow”) $T$-dependent (or propagation length $Z$-dependent) modulation of $\Re[n(x,y,T)]$ (or $\Re[n(x,y,Z])$, which is one of the management techniques stabilizing an ST soliton in BEC Kengne et al. (2021). In a mode-locked laser, it is also possible to identify a “fast” coordinate (local time $t$) associated with the frame co-moving with a soliton. The dependence of trapping potential on that coordinate allows a 3D confinement corresponding to a “pancake-like” potential in BEC. In a laser, this corresponds to an active mode-locking driven by an external synchronous phase modulation Smith et al. (1993); Wabnitz (1993); Chang et al. (2009), which varies as $\Re[\Delta n(t)]\propto\cos(2\pi t/T_{R})$, where $T_{R}$ is a cavity round-trip period. This modulation is slow (pulse duration $\tau_{p}$ is much shorter than $T_{R}$) and can be described by $\Re[\Delta n(t)]\propto t^{2}+\mathrm{H.O.T.}$. It was found that such a modulation stabilizes a driven cavity DS Leo et al. (2010); Englebert et al. (2021) and STDS Kalashnikov and Wabnitz (2022) [^1] We note that the driving (modulating) “force” can differ: synchronous modulation of complex nonlinear refraction index (i.e., phase/loss in an oscillator), synchronous pumping, and applying an external coherent pulse. One must comprehend the inherent physical differences between them..
Here, we intend to exploit a concept of 3D dissipative confinement to develop an STML mechanism to generate stable and energy-scalable STDS (“light bullet”). This technique can be considered as a road to forming a stable mass-scalable STDS in BEC and pattern formation in dissipative turbulent systems, which could be modeled by the dissipative version of the Gross-Pitaevskii equation (GPE) Carretero-González et al. (2008); Zhu et al. (2022). In contrast to Kalashnikov and Wabnitz (2020, 2022), a transversely graded dissipation (i.e., $\Im[n(x,y)]\neq 0$) will be supplemented by dissipative trapping along $t$-coordinate ($\Im[n(t)]\neq 0$), which is a result of periodic loss modulation (or synchronous pumping) synchronized with the laser cavity round-trip. For BEC, this corresponds to a “pancake-like” dissipative confining potential.
Table 1: Correspondences between photonics and BEC Kalashnikov and Wabnitz (2021)
Laser BEC
Propagation coordinate $Z$leftrightarrow time $T$
Pulse local time $t$leftrightarrow third spatial coordinate $z$
Diffraction + anomalous GDD $\leftrightarrow$ boson kinetic energy
Kerr-nonlinearity $\leftrightarrow$ boson attractive colliding potential
GRIN $\leftrightarrow$ cigar-like ($x$, $y$) trapping potential
Non-selective loss $\leftrightarrow$ homogeneous BEC dissipation
@@
Graded gain + local time modulation $\leftrightarrow$ pancake ($x,y,z$)-shaped inflow of bosons
Spectral dissipation $\leftrightarrow$ “kinetic cooling”
@@ 2.1 Model
Let us consider a waveguide laser composed of a fiber or crystalline waveguide active medium enclosed in a Fabri-Perót or ring resonator and driven by a pump beam with a parabolic-like transverse $G(x,y)-$ profile and a pump intensity varying with time $t$ at the frequency synchronized with a laser repetition rate. The $t-$coordinate corresponds to a “local time” (in a coordinate frame co-moving with an STDS) in photonics or to a longitudinal space coordinate for BEC (see Fig. 1 and Table 1) Kalashnikov and Wabnitz (2021).
Figure 1: Top: Multi-scale scheme with a local-time coordinate $t$ synchronized with an STDS group-velocity and a global evaluation coordinate $Z$ (see Table 1). The gain (filled by a red) is profiled along the radial coordinate $\sqrt{x^{2}+y^{2}}$ and the local time $t$. The last is synchronized with a laser round-trip period. Bottom: Synchronously pumped waveguide oscillator with a transversely graded refractive index. An oscillator can be linear (e.g., solid-state Cr2+:ZnS) or closed (an all-fiber laser). A pump pulse could be, for instance, an electrical pulse driving the transversely graded gain coefficient, an optical pump pulse, or an external driving pulse.
Such pump beam acts as a “soft aperture” guiding the laser beam and controlling its mode structure so that the laser field $a(z,x,y,t)$ tends to condensate into the lowest-order mode. This mechanism is akin to DKLM used to generate energy-scalable DS in solid-state lasers when a beam self-focusing enhances a spatial overlapping between pump and generation modes Zhang et al. (2022). Also, it resembles a beam-by-beam control reported in Ferraro et al. (2023). However, at variance with the latter, the beam-by-beam coupling in our case is provided not only by Kerr-nonlinearity but also spatially-profiled dissipation. Simultaneously, $t-$modulation provides confinement in the time domain. Thus, this scheme corresponds to a three-dimensional (“pancake-like”) dissipative confinement for a BEC or a hybrid mode-locking under synchronous pumping or an active assisting mode locking by synchronous loss modulation Skarka et al. (2010); Lobanov et al. (2012); Chang et al. (2009); Englebert et al. (2021). Also, $G(x,y,t)$ could be treated as one presenting a first-order expansion of a Gaussian external driving pulse.
For modeling a (3D+1)-dimensional trapped STDS, we use a distributed model describing the evolution of the $(z,x,y,t)$-dependent field $a$ propagating along the $z$-axis (see Table 1 for comparison with BEC) of an active waveguide with the linear and nonlinear refractive indexes $n(x,y)$ and $n_{2}$, respectively. Under the action of $(x,y,t)$-graded pump (Fig. 1), the evolution can be described by a variation of the driven nonlinear Helmholtz equation:
@@
where $\Delta_{x,y}$ is a Laplace operator in the Cartesian coordinates, $n_{0}$ is a waveguide cladding refractive index, $k_{0}$ is a free-space wave number, and the second-order derivative over $z$ contributes only for large beam numerical apertures NA Akhmanov et al. (1992); Demesh et al. (2023). The terms in Eq. (1) define: (I) – diffraction, (II) – group-delay dispersion with the coefficient $\beta_{2}$, (III) – spectral dissipation with the squared inverse bandwidth $\tau$, (IV) – trapping potential induced by GRIN, (V) – $(x,y,t)$- graded saturated gain with the coefficient $G$ (see footnote 1) and linear net-loss $L$, and (VI) – Kerr-nonlinearity. One has to emphasize the decisive contribution of a saturable and transversely profiled gain $G(x,y,t)$ in Eq. (1). For a low NA, one may omit the $\frac{\partial^{2}}{\partial z^{2}}$-term in Eq. (1). That leads to the nonlinear driven Schrödinger or Gross-Pitaevskii equation [^2] The terminology depends on the physical treatment of $G(x,y,t)$. It could be a gain coefficient created by synchronous pumping, active gain modulation, or a driving external pulse. In the last case, the master equation resembles the Lugiato-Lefever equation (LLE), which is a test bed model for the study of solitons and non-equilibrium patterns in driven Kerr resonators, micro-cavity and fiber lasers, VCSEL, etc. Lugiato et al. (2015); Castelli et al. (2017); Lugiato et al. (2018); Coen and Erkintalo (2016). The rescaling $\psi=a\exp(-iV_{0}z)$ ($V_{0}=(k_{0}/2n_{0})(n_{1}^{2}-n_{0}^{2})$, and $n_{1}$ is a waveguide core refractive index) leads to the dimensionless mean-field equation:
@@
$\left\{\frac{1}{2}(X^{2}+Y^{2})-\left|\psi\right|^{2}-i\left[\Lambda+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}\right]\right\}\psi,$ (2)
for the normalizations shown in Table 2 Lugiato and Lefever (1987); Raghavan and Agrawal (2000); Haelterman et al. (1992).
@@
In this work, we assume the parabolic-like transverse profile of the gain beam. Such transversely confined gain acts as a “soft aperture” with an effective size $\approx w_{p}\sqrt{G_{0}-\Lambda}$ (Table 2). Our model introduces a gain modulation along the $t-$axis described in a parabolic approximation by the modulation coefficient $\omega$ Chang et al. (2009); Lobanov et al. (2012). The last resembles the mechanism of the so-called active mode-locking Herrmann and Wilhelmi (1987). In contrast to Kalashnikov and Wabnitz (2022), the proposed method does not use additional phase modulation but only a pumping pulse train synchronized with a resonator period. For simplicity, we manipulate with a graded saturated gain $G(X,Y,t)=G(0,0,0)\times\left[1-\kappa(X^{2}+Y^{2})-\omega t^{\prime 2}\right]$, where $G(0,0,0)$ or $G_{0}$ (see Table 2) corresponds to a saturated gain along a waveguide axis at $t^{\prime}=0$, and $\kappa=G_{0}w_{p}^{-2}$, $\omega=G_{0}\mathscr{T}^{-2}$ ($\mathscr{T}$ is a “pump pulse duration”). Such an approximation is valid under the condition of quasi-steady state operation of a laser in the vicinity of the lasing threshold, where the saturated net-loss coefficient $\Lambda^{\prime}=L-G_{0}$ is negative and close to zero Kalashnikov and Wabnitz (2020).
@@
Thus, one may treat the resulting equation (2) as a dissipative extension of the driven Gross-Pitaevskii equation, which is a well-known tool for analyzing trapped BEC (Table 1) Carretero-González et al. (2008).
Table 2: Normalizations for Eq. (2)
Value Normalizations and definitions
$(X,Y)$ transverse coordinates, $(x,y)/w_{T}$
$Z$ propagation distance, $z/\zeta$
$t^{\prime}$ local time, $t/\sqrt{|\beta_{2}|\zeta}$
$\zeta$ propagation scale, $k_{0}n_{o}w_{T}^{2}$
$w_{T}$ transverse scale, $4]{w_{0}^{2}/k_{0}^{2}n_{0}\delta}$
$\delta$ GRIN contrast, $n_{1}-n_{0}$
$\left(k_{0}/2n_{0}\right)\left(n_{1}^{2}-n(x,y)^{2})\right)$ GRIN confining potential,  $\simeq k_{0}\left(n_{1}-n(r)\right)$
$\left|\psi\right|^{2}$ intensity, $k_{0}n_{2}\zeta\left|a\right|^{2}$
$\Lambda^{\prime}$ net-loss, $\Lambda-G_{0}$
$\Lambda$ averaged loss, $LL_{w}/\zeta$
$\kappa$ spatial confinement parameter, $G_{0}/w_{p}^{2}$
$\omega$ temporal confinement parameter, $G_{0}/\mathscr{T}^{2}$
$w_{p}\sqrt{G_{0}-\Lambda}$ soft aperture size
$\tau$ squared inverse spectral filter, bandwidth normalized to $|\beta_{2}|\zeta$
$k_{0}$ wave-number, $2\pi/\lambda$
$\lambda$ central wavelength
$n_{1}$ GRIN core refractive index
$n_{0}$ GRIN cladding refractive index
$n_{2}$ nonlinear refractive index
$L$ loss coefficient
$L_{w}$ waveguide (“lattice”) length
$G_{0}$ or $G(t=0,x=0,y=0)$ maximum saturated gain coefficient
$w_{p}$ pump beam size
$\mathscr{T}$ pump pulse width